OpenStudy (anonymous):

what is 2m M= 3 5 -4 0 A. 1 3 -6 -2 B. 6 5 -4 0 C. 6 10 -8 5 D. 6 10 -8 0

OpenStudy (anonymous):

These are matrix

myininaya (myininaya):

2M? Just multiply the entries of M by 2

OpenStudy (anonymous):

thanks got it

OpenStudy (anonymous):

\[2M=2\left[\begin{matrix}3 & 5 \\ -4 & 0\end{matrix}\right]=\left[\begin{matrix}2(3) & 2(5) \\ 2(-4) & 2(0)\end{matrix}\right]\]